### v3

# EMER-Ranker: Learning to Rank Emotion Descriptions in the Absence of Ground Truth

Abstract
With the recent success of large language models, Explainable Multimodal Emotion Recognition (EMER), also known as Descriptive MER (DMER), has attracted growing attention from researchers. Unlike traditional discriminative methods that rely on predefined emotion taxonomies, EMER aims to describe a person’s emotional state using free-form natural language, thereby enabling fine-grained and interpretable emotion representations. However, this free-form prediction paradigm introduces significant challenges in evaluation. Existing approaches either depend on ground-truth descriptions, which require extensive manual annotations and often fail to capture the full complexity of human emotions, or simplify the evaluation task by shifting focus from assessing descriptions to evaluating emotion labels. However, this simplification overlooks critical aspects such as emotional temporal dynamics, intensity, and uncertainty. To address these limitations, we propose EMER-Ranker, a novel evaluation strategy that reformulates the traditional “prediction-ground truth” comparison into the “prediction-prediction” comparison, eliminating the need for ground-truth descriptions. We then apply the Bradley-Terry algorithm to convert pairwise comparison outcomes into model-level rankings. Additionally, we explore the potential for automatic preference prediction and introduce EMER-Preference, the first preference dataset specifically designed for human emotions. Our work advances the field of EMER and lays the foundation for more intelligent human-computer interaction systems.
1 Introduction
Explainable Multimodal Emotion Recognition (EMER) [1], also known as Descriptive MER (DMER) [2], aims to use free-form natural language to describe emotional states. Unlike traditional discriminative methods that depend on predefined emotion taxonomies [3, 4], EMER adopts a generative paradigm, offering greater flexibility in emotion representation. This enables fine-grained and interpretable emotional expression, presenting significant opportunities for advancing emotion-intelligent human-computer interaction technologies [5]. Recent advancements in Multimodal Large Language Models (MLLMs), with their rich vocabularies and multimodal understanding capabilities, have made this task feasible. Figure 1 shows the distinctions between descriptive and discriminative MER.
Figure 1: Task comparison. Discriminative MER assigns a label from a predefined taxonomy to represent a character’s emotional state. In contrast, descriptive MER provides a more detailed description, incorporating multimodal clues, fine-grained emotion labels, and emotion intensity.
However, evaluating the quality of such open-ended descriptions remains a non-trivial task. Current evaluation strategies can be broadly classified into two categories. The first approach leverages ground-truth emotion descriptions and utilizes LLMs, such as OpenAI GPT [6], to measure the similarity between predicted and ground-truth descriptions [1]. Nevertheless, emotions are inherently linked to diverse multimodal cues, including facial expressions, (micro-)gestures, head movements, hand actions, and vocal tones [7, 8, 9]. Consequently, generating a comprehensive and accurate description of a person’s emotional state is inherently challenging. Moreover, obtaining high-quality ground-truth descriptions entails substantial annotation costs [1]. Given these challenges, researchers propose bypassing the evaluation of emotion descriptions and focusing solely on emotion labels [10]. While this approach is more cost-effective, it neglects key aspects of emotional expression present in descriptions, such as temporal dynamics, intensity, and uncertainty. Furthermore, it fails to consider the existence and reasonableness of clue-based analysis, resulting in an incomplete assessment of emotion descriptions. Figure 2 summarizes the existing evaluation strategies for EMER.
Figure 2: Existing evaluation strategies for EMER. (a) Description-based evaluation, which relies on costly and often incomplete human-annotated ground-truth descriptions; (b) Label-based evaluation, which uses cheaper emotion labels; (c) Compared to description-based methods, label-based evaluation neglects critical aspects, such as emotional temporal dynamics, intensity, uncertainty, and the existence of multimodal cues, resulting in an incomplete assessment of emotion description.
Drawing inspiration from Reinforcement Learning from Human Feedback (RLHF) [11], we transform the complex problem of designing indicators to measure semantic similarity with ground-truth descriptions into a simpler problem of learning human preferences. Specifically, we propose a novel evaluation strategy EMER-Ranker, shifting the evaluation paradigm from the “prediction-ground truth” comparison to the “prediction-prediction” comparison. In our annotation process, we provide two descriptions for a video and ask human annotators to determine which description better captures the character’s emotional state. We then employ the Bradley-Terry algorithm [12] to transform these pairwise comparison results into model-level rankings. Compared to previous methods (see Figure 2), our approach addresses the need for costly ground-truth descriptions. Furthermore, it directly assesses model performance based on descriptions, rather than relying solely on emotion labels. Therefore, this method provides a valuable complement to existing evaluation strategies for the EMER task.
Beyond the evaluation strategy, we also investigate whether the aforementioned human annotation process can be automated using MLLMs, thereby further reducing evaluation costs. Current research on LLMs-as-judges primarily focuses on pure-text or vision-language inputs [13, 14], with limited exploration in the domain of human emotions, which is characterized by multimodality, temporal dynamics, and subtle variations [15, 16]. To address this gap, we conduct a pioneering study to examine the feasibility of using MLLMs as judges for human emotions. We first construct a preference dataset, EMER-Preference. To build this dataset, we provide two emotion descriptions for each video and recruit multiple annotators to evaluate which description better aligns with the character’s emotional state. Only samples with unanimous agreement among all annotators are included in the final dataset. To the best of our knowledge, this is the first preference dataset specifically designed for human emotions. We then perform a comprehensive evaluation of various MLLMs on the preference prediction task and explore strategies to improve their consistency with human annotations. The main contributions of this paper are summarized as follows:
  • (Ranking Strategy) This paper draws inspiration from RLHF and introduces EMER-Ranker, a novel ranking strategy specifically designed for EMER. In contrast to prior approaches, our method eliminates the reliance on ground-truth descriptions but can still focus on evaluating complete emotion descriptions rather than isolated emotion labels. This work represents a significant advancement and valuable complement to existing evaluation methodologies.
  • (Automated Evaluation) This paper presents a pioneering study on automatic preference prediction for human emotions. We construct EMER-Preference, the first preference dataset in this domain, and conduct a comprehensive evaluation of multiple MLLMs on preference prediction. Our work makes a great contribution to the research on MLLMs-as-judges.
  • (Discovery) We propose several evaluation metrics and examine their relationships and effectiveness within the context of EMER. Based on our findings, we further propose multiple strategies to enhance the performance of automatic preference prediction.
2 Ranking Strategy: EMER-Ranker
This paper draws inspiration from RLHF and proposes a ranking strategy that eliminates the need for ground-truth descriptions. The core idea is to transform the traditional “prediction-ground truth” comparison into the “prediction-prediction” comparison. Section 4.4 provides an example to illustrate this ranking process in detail. Specifically, suppose there are $M$ models, and let $m_{i}$ denote one of them. Consider pairwise comparisons between model $m_{i}$ and model $m_{j}$ as an example. To ensure reliability, we conduct preference comparisons across multiple samples. For each sample $x^{k},k\in[1,K]$, we prompt model $m_{i}$ and model $m_{j}$ to generate emotion descriptions for $x^{k}$, resulting in $d^{k}_{i}$ and $d^{k}_{j}$, respectively. We then recruit multiple annotators to assign preference scores to $d^{k}_{i}$ and $d^{k}_{j}$, and their majority voted results are denoted as $p^{k}_{ij}$. Figure 3 illustrates the annotation platform layout, where description 1 and description 2 correspond to $d^{k}_{i}$ and $d^{k}_{j}$, respectively.
Figure 3: Annotation platform layout. We provide annotators a video along with two descriptions and then ask them to determine which description better reflects the character’s emotional state. They can choose from three options: description 1, description 2, or tie.
By aggregating human-annotated results across all samples $p^{k}_{ij},k\in[1,K]$, we can determine the overall performance comparison between models $m_{i}$ and $m_{j}$. For instance, if humans prefer the outputs of model $m_{i}$ more frequently than those of model $m_{j}$, we consider $m_{i}$ to outperform $m_{j}$. This preference score is then binarized into 0 or 1 and recorded in a matrix $\mathbf{W}$:
$\mathbf{W}_{ij}=\begin{cases}1,m_{i}\;\text{beats}\;m_{j}\\ 0,m_{i}\;\text{loses to}\;m_{j}\\ \end{cases}.$ (1)
Next, we apply the Bradley-Terry algorithm to infer model rankings from the preference matrix $\mathbf{W}$. In this algorithm, each model $m_{i}$ is associated with a positive parameter $\theta_{i}$, which quantifies its relative strength or advantage. The probability that model $m_{i}$ outperforms model $m_{j}$ is defined as:
$P(m_{i}\;\text{beats}\;m_{j})=\frac{\theta_{i}}{\theta_{i}+\theta_{j}}.$ (2)
Combined with the observation $\mathbf{W}$, we estimate $\theta$ by maximizing the following likelihood function:
$\max L(\theta)=\prod_{i<j}\left(\frac{\theta_{i}}{\theta_{i}+\theta_{j}}\right)^{\mathbf{W}_{ij}}\left(\frac{\theta_{j}}{\theta_{i}+\theta_{j}}\right)^{1-\mathbf{W}_{ij}}.$ (3)
For computational efficiency, we optimize the log-likelihood instead:
$\min-\sum_{i<j}\left[\mathbf{W}_{ij}\ln{\theta_{i}}+(1-\mathbf{W}_{ij})\ln{\theta_{j}}-\ln{(\theta_{i}+\theta_{j})}\right].$ (4)
Finally, we use gradient descent to estimate the advantage scores for each model. Models with higher scores are ranked higher. Section 4.4 provides an example to illustrate the above process.
3 Automatic Preference Prediction
The aforementioned ranking strategy depends on human annotations. In this section, we explore the potential of a fully automatic ranking process. To achieve this, we first construct a high-quality preference dataset tailored for human emotions, named EMER-Preference. Using this dataset, we then demonstrate the feasibility of leveraging MLLMs for automatic preference prediction.
3.1 Preference Dataset: EMER-Preference
To build the dataset, we first collect 1,368 video samples and provide two descriptions for each. Then, we recruit multiple annotators to assess these descriptions in the context of the corresponding video content and determine which one more accurately reflects the character’s emotional state. To ensure annotation quality, all annotators must pass a preliminary test assessing their alignment with the general population’s understanding of emotions. Finally, we retain only samples with unanimous agreement among all annotators, resulting in a total of 574 samples. This strict selection strategy ensures the reliability of the preference labels.
(a) w/o tie
(b) w/ tie
Figure 4: Inter-annotator consistency.
Figure 4 shows the inter-annotator consistency. As illustrated in Figure 4(a) (where ties are excluded), the inter-annotator consistency reaches approximately 70%, reflecting the upper bound of agreement among human annotators. This result also suggests that different annotators can achieve a relatively consistent consensus on this task. In Figure 4(b) (where ties are included), the average consistency drops by roughly 10%. These results suggest that ties are ambiguous and uncertain, making it difficult for annotators to reach a consensus. Therefore, in the following experiments, we report performance in both two-class (excluding ties) and three-class (including ties) conditions.
Table 1 compares our EMER-Preference with existing datasets. Our dataset integrates multimodal cues and temporal dynamics, providing a valuable resource for preference learning. Notably, this is the first preference dataset specifically designed for human emotions.
Table 1: Dataset comparison. The second column indicates the modalities used in preference annotation, where “T”, “I”, “A”, and “V” denote text, image, audio, and video, respectively. MLLM-as-a-Judge employs image sequences instead of videos; thus, we use the symbol ✔ to indicate this. As shown in the table, EMER-Preference is the only dataset that incorporates both multimodality and temporal dynamics in preference annotation, with a specific focus on human emotions.
Dataset Modality Task Description
T I A V
MTBench [17] writing, math, and knowledge
FairEval [18] 80 questions including writing, role play, etc.
PandaLM [19] law, biomedical data, etc
LLMEval [20] story generation, text summarization, dialogue, etc.
EvalBiasBench [21] 80 evaluation instances
JudgeBench [22] knowledge, reasoning, math, and coding
CALM [23] Various tasks, including commonsense reasoning, math, science QA, etc
MLLM-as-a-Judge [13]
captioning, reasoning, etc
EMER-Preference (Ours) descriptive multimodal emotion understanding
3.2 MLLM-based Solutions
Since human expresses emotions through multiple modalities, we only select mainstream open-source and closed-source MLLMs that support at least audio or video inputs. We then propose four distinct strategies to realize preference prediction.
Strategy 1 (S1).
We directly input the video along with two descriptions into MLLM and prompt it to identify the preferred description. Figure 5 illustrates the template used in S1.
Figure 5: Template for S1.
Strategy 2 (S2).
S2 decomposes the preference prediction process in S1 into two steps. Specifically, we first prompt MLLM to generate a detailed description of the video, with a particular focus on the character’s emotional state. We then treat the output from the first step as the reference and ask the model to determine which of the two candidate descriptions aligns better with this reference. Figure 6 illustrates the detailed template used in S2.
Figure 6: Template for S2 and S3.
Strategy 3 (S3).
The second step in S2 relies solely on text input, which can be implemented using an external LLM. The main reason for introducing this alternative is that the current training process of MLLMs typically begins with a pretrained LLM, followed by additional multimodal training. While this enables the model to understand other modalities, the suboptimal training strategy may harm its text processing capability. Therefore, we opt to use an external LLM for the second step.
Strategy 4 (S4).
We further divide the second step (in Figure 6) into two steps, requiring the model to perform an additional reasoning process rather than directly outputting a preference decision. Our goal is to investigate whether incorporating this intermediate reasoning step can improve performance in preference prediction. Figure 7 provides the detailed template used in S4.
Figure 7: Template for S4.
Table 2 presents a comparative analysis of different strategies, detailing the number of steps, the output content at each step, and the models employed. As illustrated in Table 2, the solution chain becomes progressively longer from S1 to S4.
Table 2: Comparison of different strategies. For each strategy, we detail the number of steps involved, the output content of each step, and the model used.
Strategy Step 1 Step 2 Step 3
output model output model output model
S1 preference MLLM
S2 description MLLM preference MLLM
S3 description MLLM preference another LLM
S4 description MLLM reasoning another LLM preference another LLM
3.3 Evaluation Metrics
Recognition Performance.
This metric measures the consistency between model predictions and human annotations. The EMER-Preference dataset contains three labels: (1) the first description is preferred; (2) the second description is preferred; (3) the descriptions are tied. As discussed in Section 3.1, ties are often ambiguous and uncertain, making it difficult for annotators to reach consensus. Therefore, we report both weighted average F1-score (WAF) and accuracy (ACC) for both two-class (excluding ties) and three-class (including ties) conditions. In the following experiments, when only one metric is reported, we default to the two-class WAF.
Flip Consistency.
In our experiments, we observe that reversing the input order sometimes leads to different results. Take Figure 8 as an example. Ideally, if we reverse the order of two descriptions, we expect the model to prefer the same description. However, in practice, models may exhibit distinct preferences. We hypothesize that this may be because their preferences are driven by the presentation order rather than the content of the descriptions. Therefore, we introduce flip consistency, a metric that measures the consistency of prediction results for forward and reverse orders. This metric evaluates the model’s robustness to position bias, ensuring reliable preference predictions.
Figure 8: Flip consistency. We require the model to prefer the same description regardless of order.
4 Results and Discussions
In this section, we first investigate the potential of automatic preference prediction, identifying strategies to enhance its performance. Subsequently, we integrate the optimal automatic preference prediction approach with EMER-Ranker for model-wise ranking. All experiments were conducted on an A100 GPU. To mitigate randomness, each experiment is repeated twice, with both average scores and standard deviations reported. For S3/S4, we adopt Qwen3-14B as the default external LLM, and additional ablation studies are provided in Appendix A.
4.1 Impact of MLLM-based Solutions
Section 3.2 presents four MLLM-based solutions, and their performance differences are compared in Figure 9. For most models, S3/S4 outperform S1/S2. The key distinction between these strategies lies in whether they leverage an external LLM. These findings underscore the advantages of integrating an external LLM, as MLLMs may exhibit limitations in language understanding. However, some exceptions also exist. For instance, Qwen2.5-VL and Qwen2.5-Omni perform better with S1/S2 than with S3/S4. This is because these models retain strong language processing capabilities due to their advanced training methodologies, allowing them to interpret complex prompts. In such cases, the longer inference chains in S3/S4 may introduce error accumulation and degrade performance. Thus, a trade-off emerges: while using an external LLM can compensate for the limited language processing abilities of some MLLMs, longer inference chains may also increase error accumulation. Consequently, the optimal solution depends on the specific model.
(a) Video-LLaVA
(b) Qwen2-Audio
(c) LLaMA-VID
(d) Chat-UniVi
(e) mPLUG-Owl
(f) VideoChat
(g) VideoChat2
(h) LLaVA-Next-Video
(i) VITA-1.5
(j) PLLAVA
(k) Qwen2.5-VL
(l) Qwen2.5-Omni
[…]
4.2 Main Results
As discussed in Section 4.1, different MLLMs exhibit preferences for distinct solutions. Table 3 summarizes the performance of each model when using its optimal strategy. These results reveal that current MLLMs demonstrate relatively weak alignment with human annotations, underscoring the challenges of automatic preference prediction. Notably, flip consistency and recognition performance show low correlation. For example, Otter [24] achieves high flip consistency but performs poorly in recognition performance. It indicates that these metrics capture distinct aspects of model behavior: recognition performance evaluates the model’s accuracy in preference prediction, whereas flip consistency ensures that predictions are based on content rather than input order. Ideally, a model should excel in both metrics. Among all models, AffectGPT [2] demonstrates the best overall performance, achieving 68.90% on the two-class WAF and 85.45% on flip consistency, highlighting its superior ability to understand human emotions.
Table 3: Main results. We report the performance of different MLLMs under their optimal strategies.
Model Optimal Strategy Two-Class Performance Three-Class Performance Flip Consistency($\uparrow$)
WAF($\uparrow$) ACC($\uparrow$) WAF($\uparrow$) ACC($\uparrow$)
Open-source MLLMs
Otter [24] S4 46.97$\pm$0.34 51.69$\pm$0.53 30.48$\pm$0.38 22.56$\pm$0.09 76.05$\pm$1.13
VideoChat2 [25] S3 48.84$\pm$1.43 48.85$\pm$1.42 44.86$\pm$0.80 42.51$\pm$0.70 43.90$\pm$1.74
Chat-UniVi [26] S3 50.26$\pm$1.24 50.27$\pm$1.24 46.32$\pm$2.13 43.55$\pm$1.92 47.74$\pm$1.05
Video-ChatGPT [27] S3 50.49$\pm$0.48 50.53$\pm$0.44 46.47$\pm$0.50 44.34$\pm$0.44 48.61$\pm$0.52
mPLUG-Owl [28] S3 50.54$\pm$0.09 50.53$\pm$0.09 46.22$\pm$0.82 43.82$\pm$0.96 47.56$\pm$1.22
Video-LLaVA [29] S3 50.69$\pm$0.55 51.78$\pm$0.27 47.52$\pm$0.29 47.74$\pm$0.00 38.59$\pm$0.26
LLaMA-VID [30] S3 53.64$\pm$0.36 53.64$\pm$0.36 48.59$\pm$0.54 45.38$\pm$0.44 43.99$\pm$1.48
VideoChat [31] S3 53.90$\pm$2.94 53.91$\pm$2.93 47.07$\pm$1.31 43.29$\pm$1.13 35.89$\pm$0.87
PLLAVA [32] S3 54.07$\pm$1.06 54.35$\pm$1.07 51.92$\pm$0.50 51.13$\pm$0.44 55.14$\pm$0.78
VITA-1.5 [33] S2 56.34$\pm$0.00 60.04$\pm$0.00 54.74$\pm$0.00 58.89$\pm$0.00 41.99$\pm$0.00
LLaVA-Next-Video [34] S3 57.05$\pm$1.61 57.46$\pm$1.51 54.78$\pm$0.83 54.27$\pm$0.61 54.09$\pm$0.44
Qwen2-Audio [35] S3 60.42$\pm$0.28 60.48$\pm$0.27 58.12$\pm$0.55 57.23$\pm$0.44 62.54$\pm$0.00
Qwen2.5-VL [36] S1 64.43$\pm$0.87 65.28$\pm$0.80 62.60$\pm$0.84 64.02$\pm$0.78 62.02$\pm$0.35
Qwen2.5-Omni [37] S2 67.21$\pm$0.00 67.32$\pm$0.00 65.29$\pm$0.00 66.03$\pm$0.00 79.09$\pm$0.00
AffectGPT [2] S4 68.90$\pm$0.52 68.92$\pm$0.53 66.16$\pm$0.53 64.98$\pm$0.35 85.45$\pm$0.96
Closed-source MLLMs
GPT-4o [6] S1 59.28$\pm$0.08 59.41$\pm$0.09 56.57$\pm$0.04 53.75$\pm$0.09 64.55$\pm$0.96
Gemini-2.0-Flash [38] S1 59.80$\pm$0.88 60.39$\pm$0.89 58.14$\pm$0.65 55.66$\pm$0.61 57.14$\pm$2.26
Gemini-2.5-Flash [39] S1 60.60$\pm$0.64 61.19$\pm$0.62 58.77$\pm$0.55 57.93$\pm$0.78 64.98$\pm$1.39
GPT-4.1 [40] S1 60.75$\pm$0.18 60.75$\pm$0.18 59.23$\pm$0.35 59.67$\pm$0.26 80.84$\pm$0.52
Gemini-1.5-Pro [41] S1 60.79$\pm$0.68 61.55$\pm$0.62 59.06$\pm$0.66 60.37$\pm$0.61 62.37$\pm$0.35
Gemini-1.5-Flash [41] S1 64.64$\pm$0.12 65.19$\pm$0.18 62.55$\pm$0.03 63.50$\pm$0.09 72.04$\pm$0.44
4.3 Model-Based Crowdsourcing
Given the limitations of MLLMs in preference prediction (see Table 3), we propose enhancing their performance through model-based crowdsourcing. Our approach consists of two key steps: 1) Filtering models based on their recognition performance and flip consistency, retaining only those that exceed 60% in both metrics. 2) Applying majority voting to aggregate predictions from the selected models. In Figure 10, we rank the filtered models by recognition performance and select the top-$N$ models for aggregation. Our experiments demonstrate that this strategy improves preference recognition results. But as the number of models $N$ increases, performance initially improves and then declines. These findings highlight the importance of selecting an appropriate $N$: a large value introduces excessive noise, while a small value limits the effectiveness of this approach.
(a) Two-Class Performance
(b) Three-Class Performance
Figure 10: Effectiveness of model-based crowdsourcing. We rank the filtered models based on their recognition performance and select the top-$N$ models for model-based crowdsourcing.
4.4 Practical Implementation: EMER-Ranker with Automatic Preference Prediction
This section illustrates our ranking process through a concrete example. We first select $M=10$ models $\mathbf{U}=${Chat-UniVi, LLaMA-VID, mPLUG-Owl, Video-ChatGPT, Otter, SALMONN, Qwen-Audio, Video-LLaVA, VideoChat, VideoChat2} and conduct preference comparisons across $N=332$ samples. This process requires a total of $C(M,2)\times N=14,940$ comparisons. To optimize evaluation efficiency, we apply the strategy described in Section 4.3 and use the top-4 performing models $\mathbf{V}=${AffectGPT, Qwen2.5-VL, Qwen2.5-Omni, Gemini-1.5-Flash} for automated preference prediction. Notably, Gemini-1.5-Flash is the only closed-source model in this group. To minimize evaluation costs, we limit automatic preference prediction to the remaining three open-source models, i.e., $\mathbf{V}=${AffectGPT, Qwen2.5-VL, Qwen2.5-Omni}. In this example, we deliberately use different MLLMs for ranking and preference prediction, i.e., $\mathbf{U}\cap\mathbf{V}=\emptyset$. Figure 11(a) summarizes the final preference prediction results. We then transform these predictions into the matrix $\mathbf{W}$, yielding the results shown in Figure 11(b). Applying the Bradley-Terry algorithm, we derive the advantage score and rank these models in descending order: Chat-UniVi, LLaMA-VID, mPLUG-Owl, Video-ChatGPT, Otter, SALMONN, Qwen-Audio, Video-LLaVA, VideoChat, and VideoChat2.
(a) Raw Results
(b) Matrix $\mathbf{W}$
Figure 11: Visualization of raw results and matrix $\mathbf{W}$. (a) We present the pairwise comparison results between all models, reporting the number of wins, losses, and ties for each model pair. (b) In this matrix, 1 indicates that model $m_{i}$ wins against model $m_{j}$, 0 indicates that model $m_{i}$ loses to model $m_{j}$, and -1 indicates that no direct comparison is performed between them.
5 Discussion: EMER-Ranker with Manual Preference Annotations
As illustrated in Figure 10, the automatic strategy achieves a maximum accuracy of only approximately 70%, highlighting its inherent limitations for preference prediction. Consequently, human annotation remains indispensable for obtaining reliable ranking results. To balance accuracy requirements with evaluation costs, we propose two practical strategies for human annotation. 1) Reduce the number of comparison samples; 2) Implement a hierarchical comparison strategy. For example, we can divide $M$ models into two subsets, conduct pairwise comparisons within each subset to identify the top-performing model from each, and then perform a final comparison between the two subset winners. This hierarchical approach significantly reduces the total number of required comparisons from the original $C(M,2)\times N$ to $C(\frac{M}{2},2)\times N+N$. Further experiments involving human annotations will be explored in our future work.
6 Conclusion
This paper proposes EMER-Ranker, a novel ranking strategy for the EMER task that eliminates the need for ground-truth descriptions. To achieve this, we shift from the traditional “prediction-ground truth” comparison paradigm into the “prediction-prediction” comparison paradigm, subsequently employing the Bradley-Terry algorithm to derive model rankings from pairwise comparison outcomes. To enhance evaluation efficiency, we explore the potential of automatic preference prediction and introduce EMER-Preference, the first preference dataset focused on human emotions. Extensive experiments across multiple MLLMs demonstrate that current models struggle with accurate preference predictions. Although model-based crowdsourcing improves performance, the predictions still show limited consistency with human annotations. Therefore, at the current stage, automatic evaluation should be treated as a reference rather than a definitive measure. For reliable ranking results, human annotation remains essential.
Appendix A Impact of External LLMs on S3/S4
As discussed in Section 3.2, S3 and S4 depend on external LLMs. In this section, we conduct a detailed investigation into the effects of different LLMs. Beyond recognition performance and flip consistency, we introduce a new metric called multi-run consistency, which measures the stability of a model’s outputs across multiple runs. As illustrated in Figure 12(a), S4 shows lower multi-run consistency than S3. These results indicate that the longer reasoning process in S4 tends to produce greater cumulative differences and more varied prediction outcomes. Figures 12(b) and 12(c) demonstrate that Qwen3-14B achieves the highest flip consistency while maintaining competitive recognition performance, highlighting its exceptional capability for preference prediction. Therefore, we adopt Qwen3-14B as the default external LLM for S3/S4.
(a) Multi-run Consistency
(b) Recognition Performance
(c) Flip Consistency
Figure 12: Impact of external LLMs. We report two-class WAF scores to evaluate recognition performance. In this table, we reveal the impact of external LLMs on AffectGPT.
Appendix B Combination of Forward and Reversed Predictions
As demonstrated in Section A, current models demonstrate suboptimal performance regarding multi-run consistency and flip consistency. This observation naturally raises the question: Can we enhance the reliability of the results by aggregating the outputs of forward and backward inputs? To evaluate this hypothesis, we execute each method twice in forward order and twice in reverse order, generating four preference descriptions $\{p_{1},\cdots,p_{4}\}$. We then employ majority voting to determine the final outcome. Figure 13 presents the impact of this strategy across different MLLMs. We observe that most models benefit from this approach, including OpenAI GPT and Gemini. However, some exceptions also exist, such as Chat-UniVi and Qwen2.5-VL, indicating that the effectiveness of this strategy is model-dependent.
(a) VideoChat2
(b) Chat-UniVi
(c) Video-ChatGPT
(d) Video-LLaVA
(e) Qwen2.5-VL
(f) AffectGPT
(g) Otter
(h) mPLUG-Owl
(i) LLaMA-VID
(j) VideoChat
(k) PLLAVA
(l) Qwen2-Audio
(m) GPT-4o
(n) GPT-4.1
(o) Gemini-1.5-Pro
(p) Gemini-1.5-Flash
(q) Gemini-2.0-Flash
(r) Gemini-2.5-Flash
Figure 13: Combination of forward and reversed predictions. We evaluate the performance of forward-only predictions versus forward-reversed combinations on two-class and three-class WAF.